INAF queue observing, 2015 Jun 21-22

Observer: F. Cusano, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Olga Kuhn on call
Instrument: LBC, MODS

Summary

At the sunset the sky is clear, humidity 30%

Night duration (12\xB0 el twilight) 20:30-4:00 ( 3:30-11:00 UT) = 7 h 30m

Observing time: 7 hrs 30m

Weather time loss: 0 hrs (0%)

Data Summary

At the sunset the sky is clear, humidity 30%.

Project_RUN Objects Filter ExpTime Notes
LBC flat field   r and i sloan    
standard star HIP56736_A0V 210_zJHK    
Piconcelli B2 1157+2724 210_zJHK 50 m  
Mignoli MOS21 SDSSJ1411 M21 red grating 1 h COMPLETED
Mignoli mos21 LBCR SDSSJ1411 M21 i sloan 30 m  
Mignoli mos21 LBCR SDSSJ1411 M21 r sloan 9 m  
Israel A4 CXOUJ204734 dual grating 1h 30m COMPLETED
Lardo M71 blue grating 1 h COMPLETED
telluric star HIP106817 210_zJHK    
Piconcelli B3 2123-0050 210_zJHK 35 m 20 m are good
Calibration LUCI        
Calibration MODS        

Details

02:58 begin the LBC flat : r and i sloan

03:36 begin with the telluric star HIP56736_A0V for Piconcelli program , target B2
  • seeing 0.7''

  • problem with the guide stars, we change it

03:41 executeLUCIscript.sh tellB2_LS_JHeI_BIS.acq luci.20150622.0006-0009

03:50 executeLUCIscript.sh tellB2_LS_210_H.spec luci.20150622.0010-0011
  • seeing 0.8''

03:52 acquisition target B2
  • executeLUCIscript.sh B2_LS_H.acq luci.20150622.0012-0015

04:05 executeLUCIscript.sh B2_LS_H.acq B2_LS_210_H.spec luci.20150622.0016-0020
  • since the presence a bad column (y=500-1018) we have changed the offset.
  • seeing 0.7''
  • 04:37 seeing 0.5''
  • 05:00 seeing 0.7''
  • 05:05 airmass 1.52, we take the last image because we are out of specific.

05:13 slew to Mignoli program in parallel mode, MOS 21
  • TO configure MODS

05:16 take test images with LUCI as required from the Support Astronomer to solve the bad column luci.20150622.0021-0022

05:28 acquire the target through the repository file mods1r.20150622.0002

05:35 execMODS mods.21.6.obs mods1r.20150622.0003-0006
  • seeing 0.8''

05:37 dofpia, /first, /redonly, /X2

05:48 send the ob 5_MOS21_i.ob
  • seeing 1.0''
  • 06:10 seeing variable 0.9''-1.5''

06:29 dofpia, /redonly, /X2

06:32 send the ob 6_MOS21_r.ob
  • seeing 1.4''
  • 06:40 seeing 1''

06:48 slew to Israel program, target A4
  • acqMODS catsj204734_ls.acq mods1r.20150622.0007-0012

06:17 execMODS catsj204734_ls.obs mods1r.20150622.0013-0016, mods1b.20150622.0002-0005
  • seeing 0.9''

08.54 slew to Lardo program, target M71
  • acqMODS M71_mos.acq mods1b.20150622.0006-0008

09:07 execMODS M71_mos.obs mods1b.20150622.0009-0011==
  • seeing 1.0''

10:15 slew to Piconcelli program

10:31 acquisition telluric star
  • executeLUCIscript.sh tellB3bis_LS_JHeI.acq luci.20150622.0030-0032
  • seeing 0.6-0.7

10:37 executeLUCIscript.sh tellB3_LS_210_H.spec luci.20150622.0033-0034

10:39 acquisition target B3
  • executeLUCIscript.sh B3_LS_H.acq luci.20150622.0035-0038

10:39 executeLUCIscript.sh B3_LS_210_H.spec luci.20150622.0039-0045

10:53 seeing jumps from 0.7'' to 2'', after few minutes it has come back to 0.7

10:55 problem with the offset of the ob, the script was wrong with absolute offset. 0039-0040 no good
  • 11:27 seeing 0.7''
  • 11:26 seeing 0.8''-0.9''
  • the last image (00.45) is taken when there is too much light in the sky

11:45 begin the calibrations flat and arc for Piconcelli B3
  • flash detector luci.20150622.0046-0047
  • executeLUCIscript.sh B3_LS_210_H.flat luci.20150622.0048-0057
  • executeLUCIscript.sh B3_LS_210_H.arc luci.20150622.0058-0061
  • arc faint, increase time exposure to 20 sec executeLUCIscript.sh B3_LS_210_H.arc luci.20150622.0063-0066
  • luci.20150622.0062 not good

12.00 calibration for Piconcelli B2
  • executeLUCIscript.sh B2_LS_210_H.arc luci.20150622.0067-0070
  • same configuration B3

12:05 calibration for Israel A4
  • execMODS catsj204734_ls_flat.cal mods1r.20150622.0017-0021, mods1b.20150622.0012-0016
  • execMODS catsj204734_ls_arc.cal mods1r.20150622.00-22, mods1b.20150622.0017

12:22 calibration for Lardo M71
  • execMODS mods.3.M71.cal mods1b.20150622.0018-22==

-- %USERSIG{RobertaCarini - 2015-06-21}%

Comments

 
Topic revision: r8 - 25 Jun 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback